OpenStudy (anonymous):

Solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (anonymous):

-y= -0.75 - 3.25x y= 0.75 + 3.25x put it in 1st equation... got it..??

OpenStudy (anonymous):

Got it